OpenStudy (anonymous):

Solve 5x^2 - 7x + 2 = 0 by completing the square. What is the constant added on to form the perfect square trinomial? a) 49/4 b) 49/25 c) 49/100

OpenStudy (primeralph):

Same story as before.

OpenStudy (anonymous):

I don't get how to do these.

OpenStudy (primeralph):

@Nurali Got this?

OpenStudy (nurali):

5x^2 - 7x + 2 Divide by 5 x^2 - 7x/5 + 2/5 = 0 x^2 - 7x/5 = -2/5 Take half of the constant for the x term and square it: half of 7/5 is 7/10, squared is 49/100. Add this to both sides of the equation x^2 - 7x/5 + 49/100 = -2/5 + 49/100 x^2 - 7x/5 + 49/100 =9/100 (x - 7/10)^2 = 9/100 i think 49/100